OpenStudy (alex_mattucci):

If d = p + 0.5at^2, which equation is solved for t?

OpenStudy (alex_mattucci):

@Michele_Laino

OpenStudy (alex_mattucci):

@pooja195

pooja195 (pooja195):

@mathmale

OpenStudy (just_one_last_goodbye):

@Alex_Mattucci it says "which equation" then there must be answer choices ^_^

OpenStudy (alex_mattucci):

yeah... forgot to put those in

OpenStudy (just_one_last_goodbye):

First of, If my memory is not failing me (all that are watching please correct)... we subtract P from both sides

OpenStudy (just_one_last_goodbye):

Second we would divide the result expression by 0.5a

OpenStudy (just_one_last_goodbye):

Then finally we figure out the square root of both sides ^_^

OpenStudy (alex_mattucci):

A. Sqrt d-p-0.5/a B. Sqrt d+p-0.5/a C. Sqrt d+p/0.5a D. Sqrt d-p/0.5a

OpenStudy (just_one_last_goodbye):

Hope that helps c:

OpenStudy (alex_mattucci):

Oh so that would be D!

OpenStudy (alex_mattucci):

A. t=Sqrt d-p-0.5/a B. t=Sqrt d+p-0.5/a C. t=Sqrt d+p/0.5a D. t=Sqrt d-p/0.5a

OpenStudy (michele_laino):

first step: if I subtract \(p\) from both sides, I get: \[d - p = p + \frac{1}{2}a{t^2} - p\] please simplify

OpenStudy (alex_mattucci):

forgot the t= in those answers above

OpenStudy (alex_mattucci):

that would be d-p=1/2at^2

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

second step: if I multiply both sides by \(2/a\), I get: \[\left( {d - p} \right)\frac{2}{a} = \frac{2}{a} \cdot \frac{a}{2}{t^2}\] please simplify

OpenStudy (alex_mattucci):

d-p/0.5a=t^2

OpenStudy (michele_laino):

hint: left side is: \[\left( {d - p} \right)\frac{2}{a}\] right side: what is: \[\frac{2}{a} \cdot \frac{a}{2} = ...?\]

OpenStudy (alex_mattucci):

i don't know

OpenStudy (michele_laino):

hint: 

OpenStudy (alex_mattucci):

oh so they cancel out

OpenStudy (michele_laino):

yes! so, what do you get?

OpenStudy (alex_mattucci):

t^2

OpenStudy (alex_mattucci):

on the right side

OpenStudy (michele_laino):

that's right! so we can write this: \[{t^2} = \left( {d - p} \right)\frac{2}{a}\]

OpenStudy (michele_laino):

finally, I take the square root of both sides, so I can write: \[\huge \sqrt {{t^2}} = \sqrt {\left( {d - p} \right)\frac{2}{a}} \] please simplify the left side

OpenStudy (alex_mattucci):

t

OpenStudy (michele_laino):

that's right! so we get: \[\huge t = \sqrt {\left( {d - p} \right)\frac{2}{a}} \] more precisely, we have: \[\Large \sqrt {{t^2}} = \pm t\] since time is positive, then we take the positive square root

OpenStudy (alex_mattucci):

ok

OpenStudy (michele_laino):

:)

OpenStudy (alex_mattucci):

so now what?

OpenStudy (michele_laino):

we have finished, you got the right answer: \[t = \sqrt {\left( {d - p} \right)\frac{2}{a}} \]

OpenStudy (alex_mattucci):

but those are not a, b, c, or d

OpenStudy (michele_laino):

we can write the right side as below: \[\Large t = \sqrt {\left( {d - p} \right)\frac{2}{a}} = \sqrt {\frac{{d - p}}{{a/2}}} = \sqrt {\frac{{d - p}}{{0.5a}}} \] since \(1/2=0.5\)

OpenStudy (alex_mattucci):

oh ok, that makes more sense!

OpenStudy (michele_laino):

:)

OpenStudy (michele_laino):

so, what is the right option?

OpenStudy (alex_mattucci):

Thank you so much for all your help! The answer is D!

OpenStudy (michele_laino):

that's right!